A digital system board uses 40 ICs manufactured with two different independent
processes. 25 ICs are manufactured using process 1 that has 80 % yield and 90 % fault
coverage. The cost of these ICs is 10 KD each. The remaining ICs (15) are manufactured
using process 2 with 85 % yield and 95 % fault coverage. The cost of these ICs is 20 KD
each. The board manufacturer spends 50 KD on the design and test of each board (even
those who are found defective). Assume all boards found defective are discarded and the
test used to detect such boards is perfect, what is the average cost of manufacturing a good
board? Hints: The cost of the defective boards need to be included in the total cost. Also,
use the DL for the unclustered random defects.
Answer #1

Cost of manufacturing the 1st 25 ICs from process 1 = 25 \times 10 KD = 250 KD

Cost of manufacturing the next 15 ICs from process 2 = 15 \times 20 KD = 300 KD

Money spent n design and test = 40 \times 50 KD = 2000 KD

Total money spent = 2000 + 250 + 300 KD = 2550 KD

Number of good ICs from process 1 = 80% of 25 = 0.8 \times 25 = 20

Number of good ICs from process 2 = 85% of 15 = 0.85 \times 15 = 12.75 \approx 12

So, cost per IC = 2550 32 = 79.69 KD